OpenStudy (anonymous):

Which is the greatest mass? Choices: 1 decigram 1 hectogram 1 milligram 1 kilogram

OpenStudy (apoorvk):

deci=0.0001 kilo hecto=0.1 kilo mili=0.000001 kilo and of course kilo is kilo! ;) now you think! and answer and verify and enjoy!!

OpenStudy (lgbasallote):

1 kilogram...deci = 10^-1 hecto = 10^2 milli = 10^-3 kilo = 10^3

OpenStudy (anonymous):

Kilo is the largest